# NFL-BA: Near-Field Light Bundle Adjustment for SLAM in Dynamic Lighting

### v3

@@ Abstract
Figure 1: NFL-BA enhances tracking and mapping in neural rendering-based SLAM (e.g., MonoGS MonoGS ) by explicitly modeling dynamic near-field lighting, with applications in endoscopy.
@@ 1 Introduction
Figure 2: We compare MonoGS performance under (1) distant static lighting and (2) dynamic near-field lighting from a co-located flashlight. Standard photometric BA performs well under static lighting but fails under dynamic lighting, degrading both trajectory and map quality. NFL-BA restores performance under dynamic lighting, matching the quality of the static-light setup.
Simultaneous Localization and Mapping (SLAM) enables autonomous agents to build a spatial map of an unknown environment while estimating their own poses within it, with wide‐ranging applications in robotics, computer vision, autonomous vehicles, and scientific imaging. Most SLAM systems sandstrom2023point ; palomer2019inspection ; chaplot2020neural ; NICESLAM2022 ; Zhu2024NICER ; NERFSLAM2022 ; yan2024gs ; 3dgs ; Matsuki_2024_CVPR ; hhuang2024photoslam assume an autonomous agent navigating an environment with distant, static illumination, e.g., a self-driving car in the streets, and they optimize a Photometric Bundle Adjustment loss where they minimize an error between the captured image and the re-rendered image using estimated 3D scene and camera poses.
However, many scientific and safety‐critical applications demand that autonomous agents operate in environments devoid of external illumination, relying instead on self-mounted light sources. For example, in endoscopy procedures, a slender flexible tube with a co-located light and camera is used to inspect internal organs such as the airway and the colon EndoDAC2024 ; Wan_EndoGSLAM_MICCAI2024 ; EndoSLAM ; liu2024endogaussian ; Yan_Deform3DGS_MICCAI2024 ; Hua_Endo4DGS_MICCAI2024 . Accurate trajectory estimation is crucial for reliably guiding instruments to areas of interest, mapping anomalies, and avoiding tissue damage during navigation. In subterranean search‐and‐rescue or collapsed‐building inspection, robots rely on onboard lamps to explore unstable voids; slight errors in pose estimation can accumulate into large drift, leading to misaligned maps, missed victims, or costly back‐tracking.
@@
In this paper, we specifically focus on demonstrating how NFL-BA can improve the performance of existing SLAM systems for 3D reconstruction and localization from endoscopy videos. SLAM can enable autonomous navigation through internal organs and guide physicians to unsurveyed regions to improve physicians’ situational awareness by providing 3D visualizations, and can help measure organ shapes, e.g. the cross-sectional area of the upper airway to diagnose airway abnormalities airway . We evaluated NFL-BA with two state-of-the-art 3DGS-based SLAM systems, general-purpose MonoGS MonoGS and endoscopy-specific EndoGSLAM Wan_EndoGSLAM_MICCAI2024 , and one neural implicit SLAM, NICE-SLAM NICESLAM2022 , by replacing their Photometric Bundle Adjustment loss with NFL-BA loss. We observe that the NFL-BA loss improves the performance of all SLAM algorithms on average when using ground-truth or estimated depth maps on the C3VD colonoscopy dataset. For example, NFL-BA significantly improves over MonoGS by reducing camera tracking error by 37% (3.48mm to 2.18mm) and camera mapping error by 38% (1.59mm to 0.99mm) when initialized by PPSNet depthPPSNet .
@@ 2 Related Works
Dense SLAM and Bundle Adjustment. Early SLAM pipelines focused on sparse feature matching for pose estimation and mapping ORBSLAM ; Campos2021ORBSLAM3 ; droid_slam ; Engel2018Direct . With advancements in neural scene representations several proposed SLAM frameworks NICESLAM2022 ; Zhu2022neuralrgbd generate dense, pixel-level that yield more detailed and robust reconstructions. More recently, 3D Gaussian surface methods have demonstrated real-time rendering with high-fidelity mapping3dgs ; Matsuki_2024_CVPR ; yan2024gs ; hhuang2024photoslam ; deng2024compact3dgaussiansplatting ; yugay2024gaussianslamphotorealisticdenseslam .
These dense SLAM approaches all rely on a core Bundle Adjustment step. Bundle Adjustment (BA) alternatively optimizes camera parameters and surface geometry by minimizing errors across multiple frames. Traditional geometric BA aligns detected 2D feature points to their 3D counterparts by minimizing reprojection error, assuming static lighting and Lambertian surfaces Hartley2003 . Although effective in controlled environments, it struggles in complex or low-texture scenes. Photometric BA (Photo-BA) alismail2016photometricbundleadjustmentvisionbased incorporates pixel intensities into the optimization process, minimizing photometric re-projection errors and proving advantageous in environments where feature matching fails Engel2018Direct . However, Photo-BA does not exploit the correspondence cues provided by dynamic or near-field lighting where image intensities vary across frames.
Near-field Lighting models. Near-field lighting has been leveraged for 3D reconstruction tasks like monocular depth and surface normal estimation PPSNet ; Zhou2009 and Photometric Stereo Lichy_2022_CVPR . Some of these approaches PPSNet ; Lichy2022 use a near-field lighting representation as input to a CNN along with captured images for predicting surface normal and geometry. In the context of Endoscopy, LightDepth rodriguez2023lightdepth and PPSNet PPSNet demonstrated the effectiveness of near-field lighting to enhance depth estimation. LightNeus batlle2023lightneus exploited the inverse-square law for light decay to improve endoscopic surface reconstruction, however with known camera parameters and pre-operative 3D CT scan.
@@
Dynamic Lighting in SLAM. Visual SLAM performance often degrades under illumination changes such as exposure shifts, specularities, and varying color temperature. Early photometric calibration methods jointly optimize camera intrinsics, exposure, and scene depths to normalize brightness variations in real time Engel2018Direct while probabilistic SLAMs with unscented filtering further stabilizes pose estimates under uncertain lighting conditions MartinezCantin2016 . More recently, learning-based matchers Sun2023LightGlue ; Zhang2023AirSLAM adapt descriptors to cope with complex lighting variations. None of these methods, however, explicitly model near‐field lighting geometry to handle this co-located light setting.
SLAM in endoscopy. Early works Stoyanov2005Soft ; EKF-mono-SLAM demonstrated the feasibility of applying SLAM in such environments by addressing dynamic lighting and tissue deformation. Researchers have often used a mixture of supervised learning on synthetic and self-supervised learning on real endoscopy datasets for tailoring SLAM frameworks to endoscopy with complex camera motion ma2019real ; zhang2021colde ; NormDepth and developed novel endoscopy SLAM frameworks rodriguez2024nr ; morlana2024topological ; iranzo2024endometric . However these techniques often struggle with challenging sequences from both synthetic and clinical data. Recently, neural rendering-based methods shi2023colonnerf ; liu2024endogaussian ; Wan_EndoGSLAM_MICCAI2024 ; Yan_Deform3DGS_MICCAI2024 ; Guo_FreeSurGS_MICCAI2024 ; Hay_Online_MICCAI2024 have proved especially effective in generating high-quality details and modeling textureless regions with a large number of Gaussians. In this work, we adopt neural rendering approaches and explicitly model the near-field lighting effects, alleviating dynamic lighting challenges and improving performance.
@@ 3 Background
To optimize $P_{t}$ and $\Theta$, dense SLAM methods typically use rendering loss $\mathcal{L}_{ren}$ , reducing the rendering errors between the rendered and captured images NICESLAM2022 ; MonoGS ; yan2024gs and, if estimated or ground truth depth maps are available, an additional depth loss $\mathcal{L}_{geo}$ can be added  tosi2024nerfs_slam_survey . Typically, these losses take the form of $L^{p}$ norm as follows with variations with $M_{t}$ as a pixel-wise mask:
@@
where $\mathcal{W}$ denotes the set of frames used for the bundle adjustment and the hyperparameters $\lambda_{ren}$ and $\lambda_{geo}$ are the loss weights. Additionally, the objective function can include any other regularization terms, such as artifact suppressing Matsuki_2024_CVPR or opacity regularization EndoGS .
@@
Implicit Neural Representations. Neural field-based SLAM methods IMAP2021 ; NICESLAM2022 ; Co-SLAM ; Zhu2024NICER ; sandstrom2023point uses a set of neural networks $F(x,d;\Theta)\rightarrow(c_{i},\sigma_{i})$, optimized to estimate the color $c_{i}$ and the volume density $\sigma_{i}$ for an input 3D coordinate $x$ and the view direction $d$. The the occupancy can be calculated from the volume density $\sigma_{i}$ and the distance between adjacent samples $\delta_{i}$ as $\alpha_{i}=1-\exp(-\sigma_{i}\delta_{i})$.
3D Gaussian Splatting. For 3D Gaussian Splatting 3dgs SLAM methods, the scene is represented by a set of Gaussians with mean $\mu^{i}$, covariance $\Sigma^{i}$ in world space, color $c_{i}$, and opacity $\alpha^{i}$. The shape parameters and occupancy $\alpha^{i}$ of the splatted 2D Gaussians are computed as follows:
@@ 4.1 Image Formation with Near-Field Lighting
Figure 4: We show that C3VD images captured with a real endoscope conform to our co-located light-camera and zero attenuation $\beta$ image formation model, as indicated by very low per-pixel scale-invariant MSE (col 3) between the original image (col 1) and the reconstructed image with masked-out specular regions (col 2).
We consider an image-formation model under near-field lighting for a single image following previous works original_pps ; PPSNet . Each pixel $p$ and the corresponding three-dimensional point $x_{p}$ in the camera space receives different light intensities and directions, characterized by the light source to surface direction $L^{d}(\cdot)$ and attenuation term $L^{a}(\cdot)$, as follows:
@@
where $\rho(\cdot)$ and $n(\cdot)$ are albedo and normal at position $x_{p}$ of pixel $p$ respectively. $PPS(\cdot)$ is a per-pixel shading term. Note that existing approaches that uses this near-field light image formation model original_pps ; PPSNet uses pixel-based representation to predict depth map or surface geometry from images captured from a single viewpoint only. In this paper, we extend the Near-Field Image Formation model beyond single-view pixel-based representation to multi-view 3D representation.
@@
Angular attenuation. Following previous works Lichy_2022_CVPR ; PPSNet , we simplify the near-field light image formation model by setting the attenuation coefficient $\beta$ in eq. 5 to zero. This effectively ignores the directional fall-off component, reducing the light attenuation term to a simple inverse-square fall-off $L^{a}(x_{p})=1/\|x_{p}-x_{L}\|^{2}$. This simplification is justified because the angular attenuation in settings like endoscopy is often negligible compared to the inverse square law attenuation, and estimating $\beta$ accurately can be challenging due to variations in endoscope designs. In future work, we plan determine the optimal value of $\beta$ for different systems and incorporate the light direction vector $r^{e}_{t}$ for more accurate modeling which can further improve camera rotation during Bundle Adjustment.
Empirical validation of our image-formation model on colonoscopy image. Fig. 4 provides an example colonoscopy image from the C3VD dataset c3vd , showing the accuracy of the near-light field model (Eq. 6) with $\beta$ of 0. Albedo was estimated by converting each RGB image to HSV color space, setting the value channel to 1 across all pixels, then converting the modified image back to RGB space. This standardizes pixel intensity variations, approximating a reflectance map where illumination effects are minimized, but does not strictly represent ground truth albedo. As shown, the image formulation model is sufficient to represent endoscopic scenes with low reconstruction errors.
@@ 4.2 Near-Field Light Bundle Adjustment Loss
Normal calculation during Bundle Adjustment. To calculate the normals $n(\cdot)$ from neural fields, we utilize the direction of the gradient of the occupancy with respect to the spatial coordinates as follows NeRD : $n(x_{i})=-\nabla\sigma(x_{i})/\|\nabla\sigma(x_{i})\|$. For Gaussian Splatting, we use the shortest axis of each Gaussian as its normal, following 10.1145/3641519.3657456 ; VCR_Gaus ; Wu2024gsrec . In both cases, we ensure the computed normal is oriented towards the camera by enforcing $n(x)^{\top}L^{d}(x)$ to be positive. Otherwise, we flip the normals by multiplying them by -1 for stability.
@@ 5 Evaluation
Table 1: Quantitative Evaluation on the C3VD c3vd dataset with oracle depth map. Replacing Photometric BA with NFL-BA significantly improves tracking quality of two state-of-the-art 3D Gaussian SLAMs, MonoGS MonoGS and EndoGS EndoGS , and one neural implicit SLAM, NICE-SLAM NICESLAM2022 .
@@
NICE-SLAM, CVPR’22 Photo 4.16 2.68 1.95
NFL 2.88 2.81 1.70
EndoGSLAM, MICCAI’24 Photo 1.93 1.81 0.85
NFL 2.04 1.13 0.97
MonoGS, CVPR’24 Photo 2.90 1.11 1.16
NFL 1.60 1.49 0.79
Our proposed method is a plug-in approach that can be applied to any existing neural-rendering-based SLAM framework. We first test our method on endoscopy videos using one neural implicit SLAM, NICE-SLAM NICESLAM2022 , as well as two existing 3DGS-SLAM frameworks: the general-purpose MonoGS Matsuki_2024_CVPR and the endoscopy-specific EndoGSLAM Wan_EndoGSLAM_MICCAI2024 . In each case, we replace the standard Photometric Bundle Adjustment loss (3) with our proposed equation NFL-BA loss (8). Additionally, we also test MonoGS Matsuki_2024_CVPR on self-captured indoor scenes with a co-located light and camera.

### v4

@@ Abstract
Figure 1: NFL-BA enhances tracking and mapping in neural rendering-based SLAM (e.g., MonoGS Matsuki et al. (2024)) by explicitly modeling dynamic near-field lighting, with applications in endoscopy.
@@ 1 Introduction
Figure 2: MonoGS performance under (1) distant static lighting and (2) dynamic near-field lighting from a co-located flashlight. Standard photometric BA performs well under static lighting but fails under dynamic lighting, degrading both trajectory and map quality. NFL-BA restores performance under dynamic lighting, matching the quality of the static-light setup.
Simultaneous Localization and Mapping (SLAM) enables autonomous agents to build a spatial map of an unknown environment while estimating their own poses within it, with wide‐ranging applications in robotics, computer vision, autonomous vehicles, and scientific imaging. Most SLAM systems Sandström et al. (2023); Palomer et al. (2019); Chaplot et al. (2020); Zhu et al. (2022, 2024); Zhi et al. (2022); Yan et al. (2024); Kerbl et al. (2023); Matsuki et al. (2024); Huang et al. (2024a) assume an autonomous agent navigating an environment with distant, static illumination, e.g., a self-driving car in the streets, and they optimize a Photometric Bundle Adjustment loss where they minimize an error between the captured image and the re-rendered image using estimated 3D scene and camera poses.
However, many scientific and safety‐critical applications demand that autonomous agents operate in environments devoid of external illumination, relying instead on self-mounted light sources. For example, in endoscopy procedures, a slender flexible tube with a co-located light and camera is used to inspect internal organs such as the airway and the colon Cui et al. (2024); Wang et al. (2024); Ozyoruk et al. (2021); Liu et al. (2024); Yang et al. (2024b); Huang et al. (2024b). Accurate trajectory estimation is crucial for reliably guiding instruments to areas of interest, mapping anomalies, and avoiding tissue damage during navigation. In subterranean search‐and‐rescue or collapsed‐building inspection, robots rely on onboard lamps to explore unstable voids; slight errors in pose estimation can accumulate into large drift, leading to misaligned maps, missed victims, or costly back‐tracking.
@@
In this paper, we specifically focus on demonstrating how NFL-BA can improve the performance of existing SLAM systems for 3D reconstruction and localization from endoscopy videos. SLAM can enable autonomous navigation through internal organs and guide physicians to unsurveyed regions to improve physicians’ situational awareness by providing 3D visualizations, and can help measure organ shapes. We evaluated NFL-BA with two state-of-the-art 3DGS-based SLAM systems, general-purpose MonoGS Matsuki et al. (2024) and endoscopy-specific EndoGSLAM Wang et al. (2024), and one neural implicit SLAM, NICE-SLAM Zhu et al. (2022), by replacing their Photometric Bundle Adjustment loss with NFL-BA loss. We observe that the NFL-BA loss improves the performance of all SLAM algorithms on average when using ground-truth or estimated depth maps on the C3VD colonoscopy dataset. For example, NFL-BA significantly improves MonoGS by reducing camera tracking error by 37% (3.48 to 2.18 mm) and camera mapping error by 38% (1.59 to 0.99 mm) when initialized by PPSNet depthParuchuri et al. (2024).
@@ 2 Related Works
Dense SLAM and Bundle Adjustment. Early SLAM pipelines focused on sparse feature matching for pose estimation and mapping Mur-Artal et al. (2015); Campos et al. (2021); Teed and Deng (2021); Engel et al. (2018). With advancements in neural scene representations several proposed SLAM frameworks Zhu et al. (2022); Azinović et al. (2022) generate dense, pixel-level that yield more detailed and robust reconstructions. More recently, 3D Gaussian surface methods have demonstrated real-time rendering with high-fidelity mappingKerbl et al. (2023); Matsuki et al. (2024); Yan et al. (2024); Huang et al. (2024a); Deng et al. (2024); Yugay et al. (2024).
These dense SLAM approaches all rely on a core Bundle Adjustment step. Bundle Adjustment (BA) alternatively optimizes camera parameters and surface geometry by minimizing errors across multiple frames. Traditional geometric BA aligns detected 2D feature points to their 3D counterparts by minimizing reprojection error, assuming static lighting and Lambertian surfaces Hartley and Zisserman (2003). Although effective in controlled environments, it struggles in complex or low-texture scenes. Photometric BA (Photo-BA) Alismail et al. (2016) incorporates pixel intensities into the optimization process, minimizing photometric re-projection errors and proving advantageous in environments where feature matching fails Engel et al. (2018). However, Photo-BA does not exploit the correspondence cues provided by dynamic or near-field lighting where image intensities vary across frames.
Near-field Lighting models. Near-field lighting has been leveraged for 3D reconstruction tasks like monocular depth and surface normal estimation Paruchuri et al. (2024); Zhou et al. (2009) and Photometric Stereo Lichy et al. (2022). Some of these approaches Paruchuri et al. (2024); Lichy et al. (2022) use a near-field lighting representation as input to a CNN along with captured images for predicting surface normal and geometry. In the context of Endoscopy, LightDepth Rodríguez-Puigvert* et al. (2023) and PPSNet Paruchuri et al. (2024) demonstrated the effectiveness of near-field lighting to enhance depth estimation. LightNeus Batlle et al. (2023) exploited the inverse-square law for light decay to improve endoscopic surface reconstruction, however with known camera parameters and pre-operative 3D CT scan.
@@
Dynamic Lighting in SLAM. Visual SLAM performance often degrades under illumination changes such as exposure shifts, specularities, and varying color temperature. Early photometric calibration methods jointly optimize camera intrinsics, exposure, and scene depths to normalize brightness variations in real time Engel et al. (2018) while probabilistic SLAMs with unscented filtering further stabilizes pose estimates under uncertain lighting conditions Martínez-Cantín and Tardós (2016). More recently, learning-based matchers Lindenberger et al. (2023); Xu et al. (2024) adapt descriptors to cope with complex lighting variations. None of these methods, however, explicitly model near‐field lighting geometry to handle this co-located light setting.
SLAM in endoscopy. Early works Stoyanov et al. (2005); Grasa et al. (2011) demonstrated the feasibility of applying SLAM in such environments by addressing dynamic lighting and tissue deformation. Researchers have often used a mixture of supervised learning on synthetic and self-supervised learning on real endoscopy datasets for tailoring SLAM frameworks to endoscopy with complex camera motion Ma et al. (2019); Zhang et al. (2021); Wang et al. (2023b) and developed novel endoscopy SLAM frameworks Rodríguez et al. (2024); Morlana et al. (2024); Iranzo et al. (2024). However these techniques often struggle with challenging sequences from both synthetic and clinical data. Recently, neural rendering-based methods Shi et al. (2023); Liu et al. (2024); Wang et al. (2024); Yang et al. (2024b); Guo et al. (2024); Hayoz et al. (2024) have proved especially effective in generating high-quality details and modeling textureless regions with a large number of Gaussians. In this work, we adopt neural rendering approaches and explicitly model the near-field lighting effects, alleviating dynamic lighting challenges and improving performance.
@@ 3 Background
To optimize $P_{t}$ and $\Theta$, dense SLAM methods typically use rendering loss $\mathcal{L}_{ren}$ , reducing the rendering errors between the rendered and captured images Zhu et al. (2022); Matsuki et al. (2024); Yan et al. (2024) and, if estimated or ground truth depth maps are available, an additional depth loss $\mathcal{L}_{geo}$ can be added  Tosi et al. (2024). Typically, these losses take the form of $L^{p}$ norm as follows with variations with $M_{t}$ as a pixel-wise mask:
@@
where $\mathcal{W}$ denotes the set of frames used for the bundle adjustment and the hyperparameters $\lambda_{ren}$ and $\lambda_{geo}$ are the loss weights. Additionally, the objective function can include any other regularization terms, such as artifact suppressing Matsuki et al. (2024) or opacity regularization Zhu et al. (2025).
@@
Implicit Neural Representations. Neural field-based SLAM methods Sucar et al. (2021); Zhu et al. (2022); Wang et al. (2023a); Zhu et al. (2024); Sandström et al. (2023) uses a set of neural networks $F(x,d;\Theta)\rightarrow(c_{i},\sigma_{i})$, optimized to estimate the color $c_{i}$ and the volume density $\sigma_{i}$ for an input 3D coordinate $x$ and the view direction $d$. The the occupancy can be calculated from the volume density $\sigma_{i}$ and the distance between adjacent samples $\delta_{i}$ as $\alpha_{i}=1-\exp(-\sigma_{i}\delta_{i})$.
3D Gaussian Splatting. For 3D Gaussian Splatting Kerbl et al. (2023) SLAM methods, the scene is represented by a set of Gaussians with mean $\mu^{i}$, covariance $\Sigma^{i}$ in world space, color $c_{i}$, and opacity $\alpha^{i}$. The shape parameters and occupancy $\alpha^{i}$ of the splatted 2D Gaussians are computed as follows:
@@ 4.1 Image Formation with Near-Field Lighting
Figure 4: Image Formation Validation. We show that C3VD images captured with a real endoscope conform to our co-located light-camera and zero attenuation $\beta$ image formation model, as indicated by very low per-pixel scale-invariant MSE between the original image and the reconstructed image with masked-out specular regions.
We consider an image-formation model under near-field lighting for a single image following previous works Iwahori et al. (1990); Paruchuri et al. (2024). Each pixel $p$ and the corresponding three-dimensional point $x_{p}$ in the camera space receives different light intensities and directions, characterized by the light source to surface direction $L^{d}(\cdot)$ and attenuation term $L^{a}(\cdot)$, as follows:
@@
where $\rho(\cdot)$ and $n(\cdot)$ are albedo and normal at position $x_{p}$ of pixel $p$ respectively. $PPS(\cdot)$ is a per-pixel shading term. Note that existing approaches that uses this near-field light image formation model Iwahori et al. (1990); Paruchuri et al. (2024) uses pixel-based representation to predict depth map or surface geometry from images captured from a single viewpoint only. In this paper, we extend the Near-Field Image Formation model beyond single-view pixel-based representation to multi-view 3D representation.
@@
Angular attenuation. Following previous works Lichy_2022_CVPR; Paruchuri et al. (2024), we simplify the near-field light image formation model by setting the attenuation coefficient $\beta$ in eq. 5 to zero. This effectively ignores the directional fall-off component, reducing the light attenuation term to a simple inverse-square fall-off $L^{a}(x_{p})=1/\|x_{p}-x_{L}\|^{2}$. This simplification is justified because the angular attenuation in settings like endoscopy is often negligible compared to the inverse square law attenuation, and estimating $\beta$ accurately can be challenging due to variations in endoscope designs. In future work, we plan determine the optimal value of $\beta$ for different systems and incorporate the light direction vector $r^{e}_{t}$ for more accurate modeling which can further improve camera rotation during Bundle Adjustment.
Empirical validation of our image-formation model on colonoscopy image. Fig. 4 provides an example colonoscopy image from the C3VD dataset Bobrow et al. (2023), showing the accuracy of the near-light field model (Eq. 6) with $\beta$ of 0. Albedo was estimated by converting each RGB image to HSV color space, setting the value channel to 1 across all pixels, then converting the modified image back to RGB space. This standardizes pixel intensity variations, approximating a reflectance map where illumination effects are minimized, but does not strictly represent ground truth albedo. As shown, the image formulation model is sufficient to represent endoscopic scenes with low reconstruction errors.
@@ 4.2 Near-Field Light Bundle Adjustment Loss
Normal calculation during Bundle Adjustment. To calculate the normals $n(\cdot)$ from neural fields, we utilize the direction of the gradient of the occupancy with respect to the spatial coordinates as follows Boss et al. (2021): $n(x_{i})=-\nabla\sigma(x_{i})/\|\nabla\sigma(x_{i})\|$. For Gaussian Splatting, we use the shortest axis of each Gaussian as its normal, following Ye et al. (2024); Chen et al. (2024); Wu et al. (2024). In both cases, we ensure the computed normal is oriented towards the camera by enforcing $n(x)^{\top}L^{d}(x)$ to be positive. Otherwise, we flip the normals by multiplying them by -1 for stability.
@@ 5 Evaluation
Table 1: Quantitative Evaluation on the C3VD Bobrow et al. (2023) dataset with oracle depth map. Replacing Photometric BA with NFL-BA significantly improves tracking quality of two state-of-the-art 3D Gaussian SLAMs, MonoGS Matsuki et al. (2024) and EndoGS Zhu et al. (2025), and one neural implicit SLAM, NICE-SLAM Zhu et al. (2022).
@@
NICE-SLAM, Photo 4.16 2.68 1.95
CVPR’22 NFL 2.88 2.81 1.70
EndoGSLAM, Photo 1.93 1.81 0.85
MICCAI’24 NFL 2.04 1.13 0.97
MonoGS, Photo 2.90 1.11 1.16
CVPR’24 NFL 1.60 1.49 0.79
Our proposed method is a plug-in approach that can be applied to any existing neural-rendering-based SLAM framework. We first test our method on endoscopy videos using one neural implicit SLAM, NICE-SLAM Zhu et al. (2022), as well as two existing 3DGS-SLAM frameworks: the general-purpose MonoGS Matsuki et al. (2024) and the endoscopy-specific EndoGSLAM Wang et al. (2024). In each case, we replace the standard Photometric Bundle Adjustment loss (3) with our proposed equation NFL-BA loss (8). Additionally, we also test MonoGS Matsuki et al. (2024) on self-captured indoor scenes with a co-located light and camera.